OpenStudy (anonymous):

express 2/5 as a decimal and how

OpenStudy (anonymous):

.4

OpenStudy (anonymous):

how do u come up with the answer 4

OpenStudy (anonymous):

point 4 or four tenths you divide the top number by the bottom number

OpenStudy (anonymous):

still don't understand

OpenStudy (anonymous):

do you have a calculator infront of you? if so type in 2 divided by 5 and you will get the point four decemil your looking for.

OpenStudy (anonymous):

4/10 now move decimal to othe side

myininaya (myininaya):

0.4 ----- 5| 2.0

myininaya (myininaya):

0.4(5)=2

myininaya (myininaya):

\[\frac{2}{5} \cdot \frac{10}{10}=\frac{20}{5 \cdot 10}=\frac{20 \cdot 1}{5 \cdot 10}=\frac{20}{5} \cdot \frac{1}{10}=4 \cdot 0.1=0.4\]

hero (hero):

myininaya working hard so...... medals